OpenStudy (anonymous):

Find f '(x) for f(x)=(x^2+1)(1+sqrt x)

OpenStudy (anonymous):

\[ f \prime = ( 2x).(1 + \sqrt{x})+(x^{2} + 1 )(\frac{ 1 }{ 2\sqrt{x} })\]

OpenStudy (precal):

product rule